OpenStudy (anonymous):

If ∠PLA and ∠ELA are complementary, ∠PLA = 5x – 2, and ∠ELA = x + 8, what is the measure of ∠PLA?

OpenStudy (anonymous):

provided these coefficients are in degrees 5x - 2 + x + 8 = 90 if they are in radians 5x - 2 + x + 8 = pi/2